OpenStudy (anonymous):

The function h is defined by h(x)=3x-2 and the function g is defined by g(x)=3h(x)-1. If g(k)=11, what is the value of k? A. 2 B. 3 C. 4 D. 5 E. 6 If k^4=16a^2, then k^2= A. 4a^2 B. 1/4a C. a/4 D. 4a E. a^2/4

OpenStudy (dtan5457):

hm

OpenStudy (dtan5457):

g(x)=3h(x)-1 h(x)=3x-2 that means g(x) is actually 3(3x-2)-1 g(x)=6x-7 x= what to make g(x)= 11?

OpenStudy (anonymous):

@dtan5457 ohhhhhh ty :D I made an error when somewhere

OpenStudy (radar):

Wouldn't that be g(x)=9x-7 ??

OpenStudy (dtan5457):

oh crud

OpenStudy (anonymous):

nvm I didn't make an error ... O.o

OpenStudy (dtan5457):

i guess i made an error too

OpenStudy (anonymous):

lol

OpenStudy (radar):

error or typo?

OpenStudy (anonymous):

nvm i got it :D

OpenStudy (dtan5457):

actual error 3(3x) literally went to 6x in my mind for some reason

OpenStudy (anonymous):

the answer is 4 ^.^

OpenStudy (dtan5457):

its been a long night ^_^

OpenStudy (anonymous):

it's been a long but short life u.u

OpenStudy (radar):

Why not 2 ??

OpenStudy (dtan5457):

9x-7 9(4)-7 36-7 if im not going crazy again is not 11

OpenStudy (anonymous):

o.0 how did u get 2?

OpenStudy (radar):

11 = 9x - 7

OpenStudy (anonymous):

3(3x-2)(x)-1=11 9x-6(x)-1=11 9x-6x-1=11 3x-1=11 3x=12 x=4

OpenStudy (radar):

maybe I erred also lol

OpenStudy (anonymous):

ye

OpenStudy (dtan5457):

this can;t be real lol

OpenStudy (anonymous):

wat can't ? @dtan5457

OpenStudy (dtan5457):

g(x)=3h(x)-1 is that 3(3x-2)-1?

OpenStudy (dtan5457):

h(x)=3x-2

OpenStudy (dtan5457):

?

OpenStudy (anonymous):

there's also another x

OpenStudy (anonymous):

3h(x)-1

OpenStudy (radar):

Would not 3h(x) = 3(3x-2) where is the extra x coming from?

OpenStudy (dtan5457):

i'm pretty sure the x is part of the function name

OpenStudy (dtan5457):

not an extra x

OpenStudy (anonymous):

function g is defined by g(x)=3h(x)-1

OpenStudy (radar):

h(x) is the function name.

OpenStudy (anonymous):

. _ . if i get it rong then whatever, but enuf w/ this qs. I'm pretty sure the other x is not part of the equation and the book this question is coming from isn't the best ._.

Nnesha (nnesha):

got it

OpenStudy (anonymous):

@Nnesha ?

Nnesha (nnesha):

3(3x-2)(x)-1=11 ^^^ how u got that x here ??

Nnesha (nnesha):

like radar sir said 11=9x-7 distribute then solve isn't that right ??